# Local EGOP for Continuous Index Learning

### v2

@@ Appendix B Proofs
In our theory, we focus on the kernel $k=\exp(-\|x\|^{2})$, as this allows for a direct correspondence to the risk distribution $\mu$ and its integration operators. We discuss general kernels in Appendix B.1.4.
@@ B.1.1 Bias Bound
To derive the upper bound from Lemma 1, we argue in stages. We first decompose the kernel smoother into two parts. The first is given by a pure Gaussian convolution and a density dependent covariance relating $f$ and $p$. Both terms can readily be controlled via the EGOP-form, leading to our bound.
@@
$=\mathbb{E}_{M}[f]+\frac{\operatorname{Cov}_{M}[f,p]}{\mathbb{E}_{M}[p]}=:\mathbb{E}_{M}[f]+A_{M}f.$
@@
This is simply the Gaussian Poincaré inequality, see for instance Bakry et al. [2013][Proposition 4.1.1]. ∎
@@
applying the Gaussian Poincaré inequality Bakry et al. [2013][Proposition 4.1.1] for the last line. ∎
@@
$\int(P_{M}-f)^{2}\,d\mu_{M}=\int(\mathbb{E}_{M}[f]-f+A_{M}f)^{2}\,d\mu_{M}=\int(\mathbb{E}_{M}[f]-f)^{2}\,d\mu_{M}+A_{M}f^{2}\leq\left(1+\frac{\operatorname{Var}_{M}(p)}{\mathbb{E}_{M}[p]^{2}}\right)W(\mu_{M}).$
Note that the density dependent factor $\frac{\operatorname{Var}_{M}(p)}{\mathbb{E}_{M}[p]^{2}}$ is bounded by the continuity of $p$ and the capped covariance assumptions.
@@ B.1.2 Variance Bound
We operate by a typical Nadaraya-Watson style argument. Our first steps are to show that the denominator of our estimator is very concentrated, allowing us to treat it as a constant up to a negligible residual. Let $\gamma_{M}\propto\det(M)^{1/2}$ by the normalization factor of $\mu_{M}$. Define
@@
$\quad\quad\leq\gamma_{M}\mathbb{E}[\exp(-(X_{i}-x)^{T}M(X_{i}-x))]=\gamma_{M}G_{M}(p).$
@@
$\mathbb{P}(|\hat{C}^{\prime}_{M}-G_{M}(p)|$ $\leq-G_{M}(p)/2)\leq 2\exp\left(-\frac{nG_{M}(p)^{2}}{8\operatorname{Var}(\gamma_{M}\exp(-(X_{i}-x)^{T}M(X_{i}-x)))+\frac{4}{3}\gamma_{M}G_{M}(p)^{2}}\right)$
$\leq 2\exp\left(-\frac{nG_{M}(p)^{2}}{8\gamma_{M}G_{M}(p)+\frac{4}{3}\gamma_{M}G_{M}(p)}\right)\leq 2\exp\left(-\frac{nG_{M}(p)}{10\gamma_{M}}\right).$
@@
$\mathbb{E}[(\hat{P}_{M}(f)-P_{M}(f))^{2}]\leq C\frac{\gamma_{M}}{n}\left(\frac{G_{2M}(p)}{G_{M}(p)^{2}}\right)+C\exp\left(-\frac{n}{\gamma_{M}}\left(\frac{G_{M}(p)}{10}\right)\right)\\ =O\left(\frac{\sqrt{\det M}}{n}\right).$
@@ B.1.3 MSE
We compile the results from the previous section to provide guarantees for the MSE and MISE against Gaussians.
@@
Compiling Lemmas 1 and 2,
@@ B.1.4 General Kernels
Substituting this into the bound above gives
@@ B.2 Recursive Kernel Learning as Variance Minimization
$\mathcal{L}(X,\lambda)=\log\det X-\lambda(\operatorname{tr}(AX)-t)\Longrightarrow\nabla_{X}\mathcal{L}=X^{-1}-\lambda A=0.$
@@ B.3.2 Supporting Arguments for Lemma 4
The first identity is definitional, the second can be found in Do Carmo [2016]. For $\|u\|=1$, we have
@@
By Lemma 14, it suffices to show that $d(\pi[x^{*}],\pi[x])=O(\|\pi_{\mathcal{T}_{x^{*}}}[x^{*}-x]\|).$ By Garcia Trillos [2019], we have that $d(\pi[x^{*}],\pi[x])=O(\|\pi[x^{*}]-\pi[x]\|)$, further reducing the problem. By our locality assumption, we can take convex combinations and remain within the tubular neighborhood, allowing us to decompose
@@
$\|D\pi_{x^{*}+t(x-x^{*})}(x-x^{*})\|=\|D\pi_{x^{*}+t(x-x^{*})}(\pi_{\mathcal{T}_{x^{*}+t(x-x^{*})}}[x-x^{*}])\|\leq C\|\pi_{\mathcal{T}_{x^{*}+t(x-x^{*})}}[x-x^{*}]\|$
$\quad\quad\leq C\|[\pi_{\mathcal{T}_{x^{*}+t(x-x^{*})}}-\pi_{\mathcal{T}_{x^{*}}}][x-x^{*}]\|+C\|\pi_{\mathcal{T}_{x^{*}}}[x-x^{*}]\|\leq Cr\|\pi_{\mathcal{T}_{x^{*}+t(x-x^{*})}}-\pi_{\mathcal{T}_{x^{*}}}\|+C\|\pi_{\mathcal{T}_{x^{*}}}[x-x^{*}]\|,$
@@ B.3.4 Taylor Expansion
where we apply Isserlis’s Theorem Isserlis [1918] to relate moments of the norm to the covariance matrix, and our capped covariance assumption is used to compress all higher order terms to quadratic order. Our proof is completed upon observing
@@
$\|\pi_{\mathcal{N}}\mathcal{L}(\mu)\pi_{\mathcal{N}}\|=\left\|\int\pi_{\mathcal{N}}\nabla f\nabla f^{T}\pi_{\mathcal{N}}\,d\mu\right\|\leq\int\|\pi_{\mathcal{N}}\nabla f\nabla f^{T}\pi_{\mathcal{N}}\|\,d\mu=\int\|[\pi_{\mathcal{N}}-\pi_{\mathcal{N}_{x}}]\nabla f(x)\nabla f(x)^{T}[\pi_{\mathcal{N}}-\pi_{\mathcal{N}_{x}}]\|\,d\mu$
@@
$\|\pi_{\mathcal{T}}\mathcal{L}(\mu)\pi_{\mathcal{N}}\|\leq\int\|\pi_{\mathcal{T}}\nabla f(x)\nabla f(x)^{T}[\pi_{\mathcal{N}}-\pi_{\mathcal{N}_{x}}]\|\,d\mu\leq C\int\|\pi_{\mathcal{T}}x\|\,d\mu\leq C\sqrt{\int\|\pi_{\mathcal{T}}x\|^{2}\,d\mu}=C\|\pi_{\mathcal{T}}\Sigma\pi_{\mathcal{T}}\|_{F}.$
@@ B.3.5 Matrix Recurrence
$B_{i+1}=t_{i+1}J(G+\beta B_{i}+(1-\beta)B_{i-1}+E_{i})^{-1}J$
where $E_{i}$ is an error term, $\beta\in(0,1)$, $G\succ 0$, $B_{0}=\alpha I$, $t_{1}=\alpha^{2}$, $B_{1}=t_{1}J(G+B_{0})^{-1}J$, $t_{i+1}=t_{i}r^{2}$, $r<1$.
@@
Suppose there exists a uniform constant $C$ such that $\|E_{i}\|\leq C\alpha$ for all $i$. Then, for $\alpha$ small enough,
$\|B_{i+1}\|=\Theta(t_{i}).$
@@
That $\alpha$ must be small is not necessary, but this will be a typical setting and will help expedite the proof. Let $\alpha<1$ be small enough that $\|B_{i}\|\leq\lambda_{\min}(G)/4$, $i=1,2$, and $C\alpha\leq\lambda_{\min}(G)/4$. We prove by strong induction that $\|B_{i+1}\|\leq t_{i+1}\frac{2}{\lambda_{\min}(G)}$. The base case is already proven, so we move on to the inductive step. We can simply compute
$G+\beta B_{i}+(1-\beta)B_{i-1}+E_{i}\succeq G/2.$
As $J$ is an orthogonal matrix, we have $\|B_{i+1}\|\leq 2t_{i+1}/\lambda_{\min}(G)$ and the claim follows.
With the decay of $B_{i}$ being settled, it follows that $B_{i}=t_{i+1}JG^{-1}J+O(t_{i+1}^{2}+\alpha)$, hence, normalizing by $t_{i+1}$, we have exact convergence to $JG^{-1}J$ in the noiseless setting. ∎
Note that the geometric decay and many of the assumptions were not completely necessary for this proof, however we introduce them now as they will greatly facilitate the following arguments.
@@
$A_{i+1}=t_{i+1}J(\beta A_{i}+(1-\beta)A_{i-1}+E_{i})^{-1}J$
where $E_{i}$ is an error term, $J$ is an involution, $\beta\in(0,1)$, $G\succ 0$, $B_{0}=\alpha I$, $t_{1}=\alpha^{2}$, $B_{1}=t_{1}J(G+B_{0})^{-1}J$, $t_{i+1}=t_{i}r^{2}$, $r<1$.
@@
Define $c(r,\beta):=\frac{1}{\beta/r+(1-\beta)/r^{2}}$ and let $X_{i}:=A_{i}/\sqrt{c(r,\beta)t_{i}}$, $R_{i}:=\sqrt{c(r,\beta)}E_{i}/\sqrt{t_{i+1}}$, and $\eta=\frac{\beta/r}{c(r,\beta)}$. Then $X_{i}$ satisfies the autonomous recurrence
$X_{i+1}=J(\eta X_{i}+(1-\eta)X_{i-1}+R_{i})^{-1}J$
@@
We work this out by simple algebra. First, as an intermediate step, introduce $C_{i}:=A_{i}/\sqrt{t_{i}}$, $R_{i}^{\prime}:=E_{i}/\sqrt{t_{i+1}}$. Then we can write
$A_{i+1}=t_{i+1}J(\beta A_{i}+(1-\beta)A_{i-1}+E_{i})^{-1}J\Longleftrightarrow\frac{A_{i+1}}{\sqrt{t_{i+1}}}=J\left(\beta\frac{A_{i}}{\sqrt{t_{i+1}}}+(1-\beta)\frac{A_{i-1}}{\sqrt{t_{i+1}}}+\frac{E_{i}}{\sqrt{t_{i+1}}}\right)^{-1}J$
$\quad\quad\Longleftrightarrow C_{i+1}=J([\beta/r]C_{i}+[(1-\beta)/r^{2}]C_{i-1}+R_{i}^{\prime})^{-1}J\Longleftrightarrow C_{i+1}=c(r,\beta)J(\eta C_{i}+(1-\eta)C_{i-1}+R_{i}^{\prime}c(r,\beta))^{-1}J$
$\quad\quad\Longleftrightarrow\frac{C_{i+1}}{\sqrt{c(r,\beta)}}=J\left(\eta\frac{C_{i}}{\sqrt{c(r,\beta)}}+(1-\eta)\frac{C_{i}}{\sqrt{c(r,\beta)}}+R_{i}^{\prime}\sqrt{c(r,\beta)}\right)J\Longleftrightarrow X_{i+1}=J(\eta X_{i}+(1-\eta)X_{i-1}+R_{i})^{-1}J.$
Thus to study $A_{i}$ it suffices to study $X_{i}$, then multiply by the corresponding factors. Our analysis proceeds as follows. First, leveraging the Thompson metric [Thompson, 1963], we show that, for summable error $R_{i}$, the spectrum of $X_{i}$ stays bounded away from 0 and $\infty$. Then, utilizing this bound, we go on to show that $X_{i}\to I$ up to some error controlled by $\alpha$. Define the Thompson metric,
@@
$e^{-L_{0}}-O(\alpha)]I\preceq X_{i}\preceq[e^{L_{0}}+O(\alpha)]I$
@@
$(1-\xi_{i})[\eta X_{i}+(1-\eta)X_{i-1}]\preceq\eta X_{i}+(1-\eta)X_{i-1}+R_{i}\preceq(1+\xi_{i})[\eta X_{i}+(1-\eta)X_{i-1}]$
for $\xi_{k}:=\|[\eta X_{i}+(1-\eta)X_{i-1}]^{-1/2}R_{i}[\eta X_{i}+(1-\eta)X_{i-1}]^{-1/2}\|.$ It follows that
$d_{T}([\eta X_{i}+(1-\eta)X_{i-1}]+R_{i},[\eta X_{i}+(1-\eta)X_{i-1}])\leq\log\frac{1+\xi_{i}}{1-\xi_{i}}$
Hence, if $e^{-L_{i}}I\preceq X_{i},X_{i-1}\preceq e^{L_{i}}I$, it follows that $\xi_{i}\leq e^{L_{i}}\|R_{i}\|$ and
$L_{i+1}\leq L_{i}+\log\frac{1+\|R_{i}\|e^{L_{i}}}{1-\|R_{i}\|e^{L_{i}}}\Longleftrightarrow e^{-L_{i+1}}\geq e^{-L_{i}}\left(\frac{e^{-L_{i}}+\|R_{i}\|}{e^{-L_{i}}-\|R_{i}\|}\right)\Longrightarrow e^{-L_{i+1}}\geq e^{-L_{i}}-2\|R_{i}\|.$
@@
Let $X_{i}$, $X_{i}^{\prime}$ be autonomous homogeneous recurrences of the Proposition 2 form, with common conjugation matrix $J$. Then
$d_{T}(X_{i+1},X_{i+1}^{\prime})\leq\max\{d_{T}(X_{i},X_{i}^{\prime}),d_{T}(X_{i-1},X_{i-1}^{\prime})\}+O(\|R_{i}\|+\|R_{i}^{\prime}\|)$
@@
Set $M_{i}=\eta X_{i}+(1-\eta)X_{i-1},M_{i}^{\prime}=\eta X_{i}^{\prime}+(1-\eta)X_{i-1}^{\prime}$. Using the invariance of the Thompson metric to conjugation by orthogonal matrices and inversion, we can compute
$d_{T}(X_{i+1},X_{i+1}^{\prime})$ $=d_{T}(\eta X_{i}+(1-\eta)X_{i-1}+R_{i},\eta X_{i}^{\prime}+(1-\eta)X_{i-1}^{\prime}+R_{i}^{\prime})$
$\leq d_{T}(M_{i},M_{i}^{\prime})+d_{T}(M_{i},M_{i}+R_{i})+d_{T}(M_{i}^{\prime},M_{i}^{\prime}+R_{i}^{\prime}).$
@@
$d_{T}(M_{i},M_{i}^{\prime})\leq\max\{d_{T}(X_{i},X_{i}^{\prime}),d_{T}(X_{i-1},X_{i-1}^{\prime})\}$
@@
for $A$ with non-vanishing spectrum. Lemma 20 verifies that the sequences $X_{i},X_{i}^{\prime}$ do not degenerate for $\alpha$ sufficiently small, thus showing the claim. ∎
@@
Let $J=I$. Then for $\alpha$ sufficiently small, $X_{i}\to I.$
@@
Define $T(x,y)=((\eta x+(1-\eta)y)^{-1},x)$, hence $T(X_{i},X_{i-1})=(X_{i+1},X_{i})$. Define
@@
hence $V(T(X_{i},X_{i-1}))\leq V(X_{i},X_{i-1})$. Thus the noiseless recurrence is autonomous, and the discrete Lasalle invariance principle [Mei and Bullo, 2017, Theorem 2] yields the existence of a set $E_{c}$, $T(E_{c})\subseteq E_{c}$ such that for all $(X,Y)\in E_{c}$, $V(E_{c})=c$ and $\lim_{i\to\infty}\inf_{E\in E_{c}}\|(X_{i},X_{i-1})-E\|\to 0$. To verify that $X_{i}\to I$ in the noiseless case, it suffices to show $E_{0}=\{(I,I)\}$ is the only invariant set with constant value function.

### v3

@@ Appendix B Proofs
In our theory, we focus on the kernel $k=\exp(-\|x\|^{2})$. We discuss general kernels in Appendix B.1.4.
@@ B.1.1 Bias Bound
To derive the upper bound from Lemma 1, we argue in stages. We first decompose the kernel smoother into two parts. The first is given by a pure Gaussian convolution and a density dependent covariance relating $f$ and $p$. Both terms can readily be controlled via the EGOP form, leading to our bound. In the following, we absorb the bandwidth into the metric and covariance, suppressing $h$ in the notation.
@@
$=\mathbb{E}_{M}[f]+\frac{\operatorname{Cov}_{M}[f,p]}{\mathbb{E}_{M}[p]}=:G_{M}[f]+A_{M}f.$
@@
This is simply the Gaussian Poincaré inequality, see for instance (Bakry et al., 2013)[Proposition 4.1.1]. ∎
@@
applying the Gaussian Poincaré inequality (Bakry et al., 2013)[Proposition 4.1.1] for the last line. ∎
@@
$\int(P_{M}-f)^{2}\,d\mu_{M}$ $=\int(\mathbb{E}_{M}[f]-f+A_{M}f)^{2}\,d\mu_{M}=\int(\mathbb{E}_{M}[f]-f)^{2}\,d\mu_{M}+A_{M}f^{2}$
$\leq\left(1+\frac{\operatorname{Var}_{M}(p)}{\mathbb{E}_{M}[p]^{2}}\right)W(\mu_{M}).$
Note that the density dependent factor $\frac{\operatorname{Var}_{M}(p)}{\mathbb{E}_{M}[p]^{2}}$ is bounded by the continuity of $p$ and the capped covariance assumptions. ∎
@@ B.1.2 Variance Bound
We operate by a typical Nadaraya-Watson style argument. Our first steps are to show that the denominator of our estimator concentrates, allowing us to treat it as a constant up to a negligible residual. Let $\gamma_{M}\propto\det(M)^{1/2}$ by the normalization factor of $\mu_{M}$. Define
@@
$\leq\gamma_{M}\mathbb{E}[\exp(-(X_{i}-x)^{T}M(X_{i}-x))]=\gamma_{M}G_{M}(p).$
@@
$\mathbb{P}(|\hat{C}^{\prime}_{M}-G_{M}(p)|$ $\leq-G_{M}(p)/2)$
$\leq 2\exp\left(-\frac{nG_{M}(p)^{2}}{8\operatorname{Var}(\gamma_{M}\exp(-(X_{i}-x)^{T}M(X_{i}-x)))+\frac{4}{3}\gamma_{M}G_{M}(p)^{2}}\right)$
$\leq 2\exp\left(-\frac{nG_{M}(p)^{2}}{8\gamma_{M}G_{M}(p)+\frac{4}{3}\gamma_{M}G_{M}(p)}\right)$
$\leq 2\exp\left(-\frac{nG_{M}(p)}{10\gamma_{M}}\right).$
@@
$\mathbb{E}[(\hat{P}_{M}(f)-P_{M}(f))^{2}]$ $\leq C\frac{\gamma_{M}}{n}\left(\frac{G_{2M}(p)}{G_{M}(p)^{2}}\right)+C\exp\left(-\frac{n}{\gamma_{M}}\left(\frac{G_{M}(p)}{10}\right)\right)$
$=O\left(\frac{\sqrt{\det M}}{n}\right).$
@@ B.1.3 MSE
We compile the results from the previous section to provide guarantees for the MISE against Gaussians.
@@
It follows immediately from Lemmas 1 and 2 that
@@ B.1.4 General Kernels
Hence,
@@ B.2 Recursive Kernel Learning as Variance Minimization
$L(X,\lambda)=\log\det X-\lambda(\operatorname{tr}(AX)-t)\Longrightarrow\nabla_{X}L=X^{-1}-\lambda A=0.$
@@ B.3.2 Supporting Arguments for Lemma 4
The first identity is definitional, the second can be found in (Do Carmo, 2016). For $\|u\|=1$, we have
@@
By Lemma 14, it suffices to show that $d(\pi[x^{*}],\pi[x])=O(\|\pi_{\mathcal{T}_{x^{*}}}[x^{*}-x]\|).$ By (Garcia Trillos, 2019), we have that $d(\pi[x^{*}],\pi[x])=O(\|\pi[x^{*}]-\pi[x]\|)$, further reducing the problem. By our locality assumption, we can take convex combinations and remain within the tubular neighborhood, allowing us to decompose
@@
$\|D\pi_{x^{*}+t(x-x^{*})}(x-x^{*})\|=\|D\pi_{x^{*}+t(x-x^{*})}(\pi_{\mathcal{T}_{x^{*}+t(x-x^{*})}}[x-x^{*}])\|$
$\quad\quad\leq C\|\pi_{\mathcal{T}_{x^{*}+t(x-x^{*})}}[x-x^{*}]\|\leq C\|[\pi_{\mathcal{T}_{x^{*}+t(x-x^{*})}}-\pi_{\mathcal{T}_{x^{*}}}][x-x^{*}]\|+C\|\pi_{\mathcal{T}_{x^{*}}}[x-x^{*}]\|$
$\quad\quad\leq Cr\|\pi_{\mathcal{T}_{x^{*}+t(x-x^{*})}}-\pi_{\mathcal{T}_{x^{*}}}\|+C\|\pi_{\mathcal{T}_{x^{*}}}[x-x^{*}]\|,$
@@ B.3.4 Taylor Expansion
where we apply Isserlis’s Theorem (Isserlis, 1918) which verifies the vanishing of third order terms, and relates the higher moments to the norm of the covariance matrix. Our capped covariance assumption is used to compress all higher order terms to quadratic order. Our proof is completed upon observing
@@
$\|\pi_{\mathcal{N}}\mathcal{L}(\mu)\pi_{\mathcal{N}}\|=\left\|\int\pi_{\mathcal{N}}\nabla f\nabla f^{T}\pi_{\mathcal{N}}\,d\mu\right\|\leq\int\|\pi_{\mathcal{N}}\nabla f\nabla f^{T}\pi_{\mathcal{N}}\|\,d\mu$
$\quad\quad=\int\|[\pi_{\mathcal{N}}-\pi_{\mathcal{N}_{x}}]\nabla f(x)\nabla f(x)^{T}[\pi_{\mathcal{N}}-\pi_{\mathcal{N}_{x}}]\|\,d\mu$
@@
$\|\pi_{\mathcal{T}}\mathcal{L}(\mu)\pi_{\mathcal{N}}\|\leq\int\|\pi_{\mathcal{T}}\nabla f(x)\nabla f(x)^{T}[\pi_{\mathcal{N}}-\pi_{\mathcal{N}_{x}}]\|\,d\mu\leq C\int\|\pi_{\mathcal{T}}x\|\,d\mu$
$\quad\quad\leq C\sqrt{\int\|\pi_{\mathcal{T}}x\|^{2}\,d\mu}=C\|\pi_{\mathcal{T}}\Sigma\pi_{\mathcal{T}}\|_{F}.$
@@ B.3.5 Matrix Recurrence
$B_{t+1}=h_{t+1}J(G+\beta B_{t}+(1-\beta)B_{t-1}+E_{t})^{-1}J$
where $E_{t}$ is an error term, $\beta\in(0,1)$, $G\succ 0$, $B_{0}=\alpha I$, $h_{1}=\alpha^{2}$, $B_{1}=h_{1}J(G+B_{0})^{-1}J$, $h_{t+1}=t_{i}r^{2}$, $r<1$.
@@
Suppose there exists a uniform constant $C$ such that $\|E_{t}\|\leq C\alpha$ for all $t$. Then, for $\alpha$ small enough,
$\|B_{t+1}\|=\Theta(h_{t}).$
@@
That $\alpha$ must be small is not necessary, but this will be a typical setting and will help expedite the proof. Let $\alpha<1$ be small enough that $\|B_{t}\|\leq\lambda_{\min}(G)/4$, $t=1,2$, and $C\alpha\leq\lambda_{\min}(G)/4$. We prove by strong induction that $\|B_{t+1}\|\leq h_{t+1}\frac{2}{\lambda_{\min}(G)}$. The base case is immediate, so we move on to the inductive step. We can simply compute
$G+\beta B_{t}+(1-\beta)B_{t-1}+E_{t}\succeq G/2.$
As $J$ is an orthogonal matrix, we have $\|B_{t+1}\|\leq 2h_{t+1}/\lambda_{\min}(G)$ and the claim follows.
With the decay of $B_{t}$ being settled, it follows that $B_{t}=h_{t+1}JG^{-1}J+O(h_{t+1}^{2}+h_{t+1}\alpha)$, hence, normalizing by $h_{t+1}$, we have exact convergence to $JG^{-1}J$ in the noiseless setting. ∎
@@
$A_{t+1}=h_{t+1}J(\beta A_{t}+(1-\beta)A_{t-1}+E_{t})^{-1}J$
where $E_{t}$ is an error term, $J$ is an involution, $\beta\in(0,1)$, $G\succ 0$, $B_{0}=\alpha I$, $h_{1}=\alpha^{2}$, $B_{1}=h_{1}J(G+B_{0})^{-1}J$, $h_{t+1}=h_{t}r^{2}$, $r<1$.
@@
Define $c(r,\beta):=\frac{1}{\beta/r+(1-\beta)/r^{2}}$ and let $X_{t}:=A_{t}/\sqrt{c(r,\beta)h_{t}}$, $R_{t}:=\sqrt{c(r,\beta)}E_{t}/\sqrt{h_{t+1}}$, and $\eta=\frac{\beta/r}{c(r,\beta)}$. Then $X_{t}$ satisfies the autonomous recurrence
$X_{t+1}=J(\eta X_{t}+(1-\eta)X_{t-1}+R_{t})^{-1}J$
@@
We work this out by simple algebra. First, as an intermediate step, introduce $C_{t}:=A_{t}/\sqrt{h_{t}}$, $R_{t}^{\prime}:=E_{t}/\sqrt{h_{t+1}}$. Then we can write
$A_{t+1}=h_{t+1}J(\beta A_{t}+(1-\beta)A_{t-1}+E_{i})^{-1}J$
$\quad\quad\Longleftrightarrow\frac{A_{t+1}}{\sqrt{h_{t+1}}}=J\left(\beta\frac{A_{t}}{\sqrt{h_{t+1}}}+(1-\beta)\frac{A_{t-1}}{\sqrt{h_{t+1}}}+\frac{E_{t}}{\sqrt{h_{t+1}}}\right)^{-1}J$
$\quad\quad\Longleftrightarrow C_{t+1}=J([\beta/r]C_{t}+[(1-\beta)/r^{2}]C_{t-1}+R_{t}^{\prime})^{-1}J$
$\quad\quad\Longleftrightarrow C_{t+1}=c(r,\beta)J(\eta C_{t}+(1-\eta)C_{t-1}+R_{t}^{\prime}c(r,\beta))^{-1}J$
$\quad\quad\Longleftrightarrow\frac{C_{t+1}}{\sqrt{c(r,\beta)}}=J\left(\eta\frac{C_{t}}{\sqrt{c(r,\beta)}}+(1-\eta)\frac{C_{t}}{\sqrt{c(r,\beta)}}+R_{t}^{\prime}\sqrt{c(r,\beta)}\right)J$
$\quad\quad\Longleftrightarrow X_{t+1}=J(\eta X_{t}+(1-\eta)X_{t-1}+R_{t})^{-1}J.$
Thus to study $A_{t}$ it suffices to study $X_{t}$, then multiply by the corresponding factors. Our analysis proceeds as follows. First, leveraging the Thompson metric (Thompson, 1963), we show that, for summable error $R_{t}$, the spectrum of $X_{t}$ stays bounded away from 0 and $\infty$. Then, utilizing this bound, we go on to show convergence of $X_{t}$. Define the Thompson metric,
@@
$e^{-L_{0}}-O(\alpha)]I\preceq X_{t}\preceq[e^{L_{0}}+O(\alpha)]I$
@@
$(1-\xi_{t})[\eta X_{t}+(1-\eta)X_{t-1}]\preceq\eta X_{t}+(1-\eta)X_{t-1}+R_{t}\preceq(1+\xi_{t})[\eta X_{t}+(1-\eta)X_{t-1}]$
for $\xi_{t}:=\|[\eta X_{t}+(1-\eta)X_{t-1}]^{-1/2}R_{t}[\eta X_{t}+(1-\eta)X_{t-1}]^{-1/2}\|.$ It follows that
$d_{T}([\eta X_{t}+(1-\eta)X_{t-1}]+R_{t},[\eta X_{t}+(1-\eta)X_{t-1}])\leq\log\frac{1+\xi_{t}}{1-\xi_{t}}$
Hence, if $e^{-L_{t}}I\preceq X_{t},X_{t-1}\preceq e^{L_{t}}I$, it follows that $\xi_{t}\leq e^{L_{t}}\|R_{t}\|$ and
$L_{t+1}\leq L_{t}+\log\frac{1+\|R_{t}\|e^{L_{t}}}{1-\|R_{t}\|e^{L_{t}}}\Longleftrightarrow e^{-L_{t+1}}\geq e^{-L_{t}}\left(\frac{e^{-L_{t}}-\|R_{t}\|}{e^{-L_{t}}+\|R_{t}\|}\right)\Longrightarrow e^{-L_{t+1}}\geq e^{-L_{t}}-2\|R_{t}\|.$
@@
Let $X_{t}$, $X_{t}^{\prime}$ be autonomous homogeneous recurrences of the form of Proposition 2, with common conjugation matrix $J$. Then
$d_{T}(X_{t+1},X_{t+1}^{\prime})\leq\max\{d_{T}(X_{t},X_{t}^{\prime}),d_{T}(X_{t-1},X_{t-1}^{\prime})\}+O(\|R_{t}\|+\|R_{t}^{\prime}\|)$
@@
Set $M_{t}=\eta X_{t}+(1-\eta)X_{t-1},M_{t}^{\prime}=\eta X_{t}^{\prime}+(1-\eta)X_{t-1}^{\prime}$. Using the invariance of the Thompson metric to conjugate by orthogonal matrices and invert, we can compute
$d_{T}(X_{t+1},X_{t+1}^{\prime})$ $=d_{T}(\eta X_{t}+(1-\eta)X_{t-1}+R_{t},\eta X_{t}^{\prime}+(1-\eta)X_{t-1}^{\prime}+R_{t}^{\prime})$
$\leq d_{T}(M_{t},M_{t}^{\prime})+d_{T}(M_{t},M_{t}+R_{t})+d_{T}(M_{t}^{\prime},M_{t}^{\prime}+R_{t}^{\prime}).$
@@
$d_{T}(M_{t},M_{t}^{\prime})\leq\max\{d_{T}(X_{t},X_{t}^{\prime}),d_{T}(X_{t-1},X_{t-1}^{\prime})\}$
@@
for $A$ with non-vanishing spectrum. Lemma 20 verifies that the sequences $X_{t},X_{t}^{\prime}$ do not degenerate for $\alpha$ sufficiently small, thus showing the claim. ∎
@@
Let $J=I$. Then for $\alpha$ sufficiently small, $X_{t}\to I.$
@@
Define $T(x,y)=((\eta x+(1-\eta)y)^{-1},x)$, hence $T(X_{t},X_{t-1})=(X_{t+1},X_{t})$. Define
@@
hence $V(T(X_{t},X_{t-1}))\leq V(X_{t},X_{t-1})$. Thus the noiseless recurrence is autonomous, and the discrete Lasalle invariance principle Mei and Bullo (2017, Theorem 2) yields the existence of a set $E_{c}$, $T(E_{c})\subseteq E_{c}$ such that for all $(X,Y)\in E_{c}$, $V(E_{c})=c$ and $\lim_{t\to\infty}\inf_{E\in E_{c}}\|(X_{t},X_{t-1})-E\|\to 0$. To verify that $X_{t}\to I$ in the noiseless case, it suffices to show $E_{0}=\{(I,I)\}$ is the only invariant set with constant value function.